OpenStudy (luigi0210):

Determine:

OpenStudy (luigi0210):

\[\int\limits sinx dx\]

OpenStudy (dls):

-cosx+c?

OpenStudy (psymon):

\[\int\limits_{}^{}\sin=666\]

OpenStudy (luigi0210):

Thanks :3

OpenStudy (dls):

\[\LARGE \frac{d}{dx}(cosx)=-sinx\] \[\LARGE \int\limits_{}^{}\sin x dx=-\cos x+c\]